OpenStudy (tanabugg):

7. A rectangular prism has a width of 92 ft and a volume of 240 ft3. Find the volume of a similar prism with a width of 46 ft. Round to the nearest tenth, if necessary. (1 point) 30 ft3 40 ft3 60 ft3 120 ft3

OpenStudy (tanabugg):

@soprano.h.d0816

OpenStudy (soprano.h.d0816):

Do you know how to find volume?

OpenStudy (tanabugg):

no

OpenStudy (soprano.h.d0816):

The cube of the scale factor of two similar rectangular solids is equal to the ratio of the volumes of the two similar solids. The scale factor in this case is the ratio of the widths: 92/23. Following the theorem, (92/23) ^ 3 = 240/V where V is the volume of the rectangular solid with width 23.

OpenStudy (tanabugg):

Would it be 60? @soprano.h.d0816

OpenStudy (soprano.h.d0816):

@agent0smith can you see if this is correct. Like, double check me lol

OpenStudy (tanabugg):

? @agent0smith

OpenStudy (tanabugg):

So I'm assuming I was right?

OpenStudy (soprano.h.d0816):

I'm saying you are correct. I'm guessing he's busy. Lol

OpenStudy (tanabugg):

Ok. Can you help me with another?

OpenStudy (soprano.h.d0816):

Sure!

OpenStudy (tanabugg):

Thanks! Hold on a sec

OpenStudy (soprano.h.d0816):

Okay(:

OpenStudy (tanabugg):

Find the value of the similar figures

OpenStudy (tanabugg):

I think it's 2cm

OpenStudy (soprano.h.d0816):

How'd you get 2?

OpenStudy (tanabugg):

The answer choices are 3 cm 18 cm 16 cm 2 cm

OpenStudy (tanabugg):

Oh I don't have my paper that I wrote my answer and work for bc my mum threw it away

OpenStudy (soprano.h.d0816):

Okay. So, what you would do is set it in a proportion. Correct?

OpenStudy (tanabugg):

Um yes

OpenStudy (soprano.h.d0816):

So, the propertion would look like: \[\frac{ 15 }{ 6 } = \frac{ 5 }{ x }\] Then you would cross multiply.

OpenStudy (tanabugg):

Yes

OpenStudy (soprano.h.d0816):

What do you get when you cross multiply? Like, what's the equation you got?

OpenStudy (tanabugg):

OH! Wait I got it I know that it's 2

OpenStudy (soprano.h.d0816):

Good!!